A corporation can earn 7.50% if it invests in municipal bonds. The corporation can also earn 8.00% (before-tax) by investing in preferred stock. Assume that the two investments have equal risk. What is the break-even corporate tax rate that makes the corporation indifferent between the two investments? Assume a 70.00% dividend exclusion for tax on dividends. (Do not round your intermediate answer and round your final answer to two decimal places.)

Group of answer choices

17.29%

26.04%

17.08%

20.83%

21.88%
Answer #1
Assume tax rate is X
municipal bonds is tax free
and 70% dividend is tax free
hence equation would be as follows for indifference tax rate
7.5 = (8*70%) + [8*30%*(1-X)]
7.5 =5.6+2.4-2.4X
7.5 =8-2.4X
2.4X =8-7.5
2.4X =0.50
X =0.20833
X =20.833%
Therefore the correct option is FOURTH